The activation energy of a certain reaction is 30.1 kJ/mol . At 28 ∘C , the rate constant is 0.0120s−1 . At what temperature in degrees Celsius would this reaction go twice as fast?
Answer #1

Given:

T1 = 28 oC

=(28+273)K

= 301 K

K1 = 1.2*10^-2 s-1

K2 = 2.4*10^-2 s-1

Ea = 30.1 KJ/mol

= 30100 J/mol

use:

ln(K2/K1) = (Ea/R)*(1/T1 - 1/T2)

ln(2.4*10^-2/1.2*10^-2) = (30100.0/8.314)*(1/301 - 1/T2)

0.6931 = 3620.3993*(1/301 - 1/T2)

T2 = 319 K

= (319-273) oC

= 46 oC

Answer: 46 oC